# Forecasting the spatiotemporal evolution of fluid-induced microearthquakes with deep learning

### v1

@@ 2.1 EGS Collab hydraulic stimulation datasets
Figure 3 presents the series of stimulations along with the spatiotemporal MEQ data and corresponding magnitudes. Detailed information about the MEQs—including location, time, and magnitudes—was continuously monitored during the hydraulic stimulations [60, 17]. In addition, to quantify the spatial extent of MEQs in response to fluid injection, we extracted the 95th and 50th percentiles (median) distances of the MEQ clouds from the injection points as a function of time. Although the monitoring array is extensive, the catalog still carries intrinsic uncertainties: hypocenter locations are accurate to about 1 m and there is no reported uncertainty range for magnitude [60]. These uncertainties limit the fidelity of the training data and establish a floor on achievable forecast accuracy.
@@ 2.2 Forecasting performance
Figure 5: Cumulative MEQ counts: observed data (black dotted) versus forecasts for the 1-second (blue), 15-second (red), and 30-second (green) models. Each forecast curve is constructed by predicting successive, non-overlapping segments whose length equals the forecast interval and concatenating them to cover the full record. Panels show the training (left), validation (middle), and test (right) sets. Shaded bands denote ±$\pm\sigma$ (one predicted standard deviation, corresponding to $\approx 68\%$ coverage under a Gaussian assumption).
@@
Figure 8: Spatial view of MEQ evolution. Rows correspond to the training (top), validation (middle), and test (bottom) datasets; each row shows projections on the XY, YZ, and ZX planes. In every projection the solid circle marks the observed 50th-percentile radius ($P_{50}$), and the dashed circle the 95th-percentile radius ($P_{95}$). Forecasts from the 1-, 15-, and 30-second models appear in blue, red, and green, respectively, with shaded rings denoting ±$\pm\sigma$ (one predicted standard deviation).
@@ 3.1 Estimation of permeability enhancement
During the first stimulation, the observed cumulative logarithmic seismic moment reaches $\approx 3$ (Figure 6 left), implying a permeability increase of roughly two orders of magnitude. The 1-second forecast reproduces this estimate, whereas the 15-second forecast model overpredicts the moment by about one order of magnitude, and the 30-second forecast model underpredicts it by a similar amount. Because the cumulative seismic moment predicted by our network can be mapped directly to permeability changes, the model provides a practical, indirect means of tracking permeability evolution during hydraulic stimulation—though this mapping is valid only for the initial seismic-moment range where the derivation’s assumptions hold.
@@ 4.1 Data preprocessing: crop and normalization
where the monitoring dimension $M=6$ includes hydraulic stimulation features —(1) flow rate ($x_{1}$) and (2) well head pressure ($x_{2}$)— and spatiotemporal MEQ features —(3) cumulative MEQ numbers ($x_{3}$), (4) $\log{M_{0}}$ ($x_{4}$), (5) 95th percentile distance ($x_{5}$), (6) 50th percentile distance)($x_{6}$).
@@ 4.3 Neural-network hyper-parameter tuning
Target features $D$ 4 per timestep
Multi-head attention 4 heads, key_dim $=M$
Feed-forward width 32 units (ReLU)
@@
Maximum epochs 100 (early stopping)
Loss core Gaussian NLL $+$ monotonic penalty

### v2

# Deep learning forecasts the spatiotemporal evolution of fluid-induced microearthquakes

@@ 2.1 EGS Collab hydraulic stimulation datasets
Figure 3 presents the series of stimulations along with the spatiotemporal MEQ data and corresponding magnitudes. Detailed information about the MEQs—including location, time, and magnitudes—was continuously monitored during the hydraulic stimulations [60, 17]. In addition, to quantify the spatial extent of MEQs in response to fluid injection, we extracted the 95th and 50th percentiles (median) distances of the MEQ clouds from the injection points as a function of time. Although the monitoring array is extensive, the catalog still carries intrinsic uncertainties: hypocenter locations are accurate to about 1 m and there is no reported uncertainty range for magnitude [60]. These uncertainties limit the fidelity of the training data and establish a floor on achievable forecast accuracy. Additionally, including all raw events—without excluding those below the magnitude of completeness—could constrain the neural network’s capability to learn underlying MEQ patterns (Fig. S1).
@@ 2.2 Forecasting performance
[…]
@@
Figure 8: Spatial evolution of microearthquake (MEQ) clouds and forecast performance. The figure shows the spatial distribution of MEQs and forecast results for different datasets and time horizons. Each row represents a dataset: training (top), validation (middle), and test (bottom). Each column shows projections on the XY, YZ, and ZX planes. Solid circles indicate the observed 50th-percentile radius ($P_{50}$), while dashed circles represent the 95th-percentile radius ($P_{95}$). Blue and red lines show forecasts from the 1-second and 15-second models, respectively, with shaded regions denoting ±$\pm\sigma$ (one predicted standard deviation). Colored dots represent MEQs from different stimulation phases: yellow for Stim 3, green for Stim 4, and purple for Stim 5. Solid black lines indicate the injection well, dashed black lines the production well, and red triangles mark the injection point. All spatial dimensions are in meters.
@@ 3.1 Estimation of permeability enhancement
During the first stimulation, the observed cumulative logarithmic seismic moment reaches 3$\approx$ 3≈ 3 (Figure 6 left), implying a permeability increase of roughly two orders of magnitude. The 1-second forecast reproduces this estimate, whereas the 15-second forecast model overpredicts the moment by about one order of magnitude, and the 30-second forecast model underpredicts it by a similar amount. Because the cumulative seismic moment predicted by our network can be mapped directly to permeability changes, the model provides a practical, indirect means of tracking permeability evolution during hydraulic stimulation—though this mapping is valid only for the initial seismic-moment range where the derivation’s assumptions hold.
@@ 4.1 Data preprocessing: crop and normalization
[…]
@@ 4.3 Neural-network hyper-parameter tuning
Forecasting features $F$ 4 per timestep
Number of multi-head attention 4 heads
Feed-forward width 32 units (ReLU activation)
@@
Maximum epochs 100 (with early stopping)